# Range Membership Inference Attacks

### v1

@@ Abstract
Machine learning models can leak private information about their training data, but the standard methods to measure this risk, based on membership inference attacks (MIAs), have a major limitation. They only check if a given data point exactly matches a training point, neglecting the potential of similar or partially overlapping data revealing the same private information. To address this issue, we introduce the class of range membership inference attacks (RaMIAs), testing if the model was trained on any data in a specified range (defined based on the semantics of privacy). We formulate the RaMIAs game and design a principled statistical test for its complex hypotheses. We show that RaMIAs can capture privacy loss more accurately and comprehensively than MIAs on various types of data, such as tabular, image, and language. RaMIA paves the way for a more comprehensive and meaningful privacy auditing of machine learning algorithms.
@@ 1 Introduction
To quantify the privacy risk of machine learning models, a privacy notion needs to be fixed first. The reigning privacy notion is defined by membership information. Membership information of a data point is binary, but this single bit of information carries huge privacy implications. Being able to infer membership information opens up the possibility of conducting data reconstruction attack [36, 17, 4, 29], where the reconstruction attack inspects the membership of plausible data points to recover the training set. The de-facto way to audit the privacy risk according to this privacy notion is to conduct membership inference attacks (MIAs) [39], where an adversary aims to predict whether a given query data is part of the training set of the target model. The more powerful the membership inference attack is, the higher the privacy risk the target model bears.
@@
Membership inference attacks provide a lower bound of the true privacy risk of a model, so improving the attack performance also means tightening the bound of privacy risk estimation. So far, the community has put much effort into improving the power of membership inference attacks by crafting better membership signals and constructing better statistical tests [38, 39, 35, 44, 5, 47]. While these have been useful for the betterment of privacy auditing, they have ignored the fundamental drawback of membership inference attacks as a practical privacy auditing tool, i.e., MIAs assume it is a privacy concern only if the adversary can identify the exact, full version of the training data. However, if the adversary can identify data points that are similar enough to the training points, it should also be treated as a significant privacy risk, because those points can contain similar levels of private information. For example, two Alice’s photos taken from slightly different angles, or with a different background would contain similar private information about Alice’s face or location. Similar to images, small perturbations or rephrasing in textual data also affect little in the sensitivity of the information conveyed [13]. This oversight means private information leakage beyond exact matches of training data is ignored and current privacy auditing tools might produce overly optimistic results.
@@
Besides, focusing on exact membership inference attacks renders them incapable of handling queries with missing values. This is another major limitation of MIAs, as data records with the same sensitive features and a few missing non-private features would carry a similar level of private information as the full data records. Imagine the case where an adversary can infer that an Asian person of age 25, identification number 123456 is in the hospital training set, there is no need to identify the rest of the features because the adversary is already able to pinpoint who has HIV from the given subset of key features. Even if the key identifiers are missing and unknown, it is still a grave privacy threat if the attacker can infer the rest of the features that often contain quasi-identifiers, which has been studied extensively in the failure of k-anonymity [37], where the attacker is able to reconstruct and identify training data with certain features removed. However, if we make up the missing values or pass noisy data to the membership inference attack, the attack is expected to output “not a member," since the chance of us coming up with the right values is so slimSupposeIf we use property inference [40, 1, 8, 9, 45, 20] to help fill in the missing features, the imputed data may be have inflated membership score even if they are non-members, leading to a higher false positive rate. Therefore, existing frameworks struggle with quantifying privacy leakage with missing features.
@@
We argue that privacy quantification should not be point-based, because a small neighborhood around training points also contains similar private information. Hence, in this paper, we are proposing a new attack framework called range membership inference attacks (RaMIAs) to better capture the notion of privacy. Instead of using point queries and testing for exact matches, range membership inference attacks use range queries that cover a set of points. The goal of range membership inference attacks is to infer if the given range query contains any training point.
@@
Range membership inference attacks extend the formulation of membership inference attacks. We adapt the original inference game formulation to reflect the change to range queries in RaMIAs. This extended formulation produces composite hypotheses in the likelihood ratio tests, which are the standard and best attack techniques in MIAs [38, 44, 5, 47]. Our method is based on standard statistical methods for composite hypothesis testing, namely generalized likelihood ratio tests and Bayes factors. We show that RaMIAs can provide a more comprehensive notion of privacy by detecting private information leakage from the vicinity of training data when MIAs underestimate such privacy risk. Specifically, we observe a simple flipping can cause the membership score to decrease from very high to 0 (Figure 1), and the overall AUC can drop 0.20 if we test image classifiers with horizontally flipped images (Figure 2(c)). RaMIA, implemented with our simple attack strategy (Sec 4), supersedes MIA by at least 5% on image datasets (Fig 3(b), 3(c)), providing better privacy auditing at the cost of as few as 15 samples, which is insignificant compared to the dimensionality of the data space.
@@
In this paper, we emphasize the motivation and formulation of our newly proposed attack framework, RaMIA. As a proof-of-concept, we experiment RaMIA with a simple attack strategy on tabular, image and text datasets, where RaMIA unanimously outperforms MIA. Additionally, our attack can also be potentially used in the pioneering membership inference and data extraction attacks on generative models [4, 43, 7], where the current evaluation requires finding the closest training image for all candidate data and computing their distances. By setting the distance function as the range function and conducting RaMIAs, we can evaluate the attacks more systematically.
@@
Figure 1: Examples of similar data with drastically different MIA scores. MIA would classify one of them as member, but the other as non-member.
@@ 2.1 Membership inference attacks
The membership inference attack (MIA) [39] is a type of inference attack against machine learning models to infer whether a given data sample is part of the model’s training set. Mathematically, given a model $f$ and a query point $\mathbf{x}$, the MIA aims to output 1 if $\mathbf{x}$ is a training point, and 0 otherwise. There are various methods to construct and conduct the attack, and it is still an active research direction that sees more powerful attacks being developed. Shokri et al. [39] use a shadow model based approach where shadow models are trained on known training sets in similar ways to the target model. Confidence values of the training and test data on the shadow models are computed, which are then used as benchmarks in testing. However, the high cost and strong assumption of knowing the target model’s training details make the attack often infeasible. Yeom et al. [45] use model loss as a signal and threshold it, scraping the need for shadow models. Then MIA is formulated as an inference game (See Sec 3.1.1). Researchers turn to the principled approach to solve the game via likelihood ratio tests [38, 5, 44, 47]. Carlini et al. [5] and Ye et al. [44] propose reference-model based approaches, where target signals are compared to those obtained on reference models to obtain the likelihood ratio. To further boost the attack power, Zarifzadeh et al. [47] assume the attacker has access to a pool of population data so that the likelihood ratio from reference-based attacks can be calibrated on ratios obtained on (non-member) population data.
Recent attacks [4, 47] further boost the attack performance on image data by augmenting the test queries with train-time augmentations. This assumes that the attacker knows the exact train-test augmentations in advance, and is able to sample from them. Augmenting training images with non train-time augmentations is not considered for valid reasons: those augmented images would be non-members in the current privacy notion.
@@ 2.2 Range queries
If we make a connection to the field of databases, a membership inference attack operates on point queries or exact match queries. That is, each query to the membership inference attack only contains one data point and the attack only concerns if this very point is in the training set. On the other hand, range query, which is also a common querying operation in database systems, wants to retrieve all data points that fall into the "range". The most fundamental difference to point query is that the retrieved result often contains multiple data points instead of a single one. Our proposed attack, the range membership inference attack, operates with range queries.
@@ 3.1.2 Evaluation of MIA
Evaluation is done with a set of training and test points. True positive rate (TPR) and false positive rate (FPR) are computed by sweeping over all possible threshold values. By plotting the receiver operating characteristic curve (ROC), the power of an attack strategy can be represented by the area under the curve (AUC). A clueless adversary who can only randomly guess the membership labels will get an AUC of 0.5. For stronger adversaries, they predict membership more correctly at each error level. Hence, they would achieve higher TPR at each FPR, and get a higher AUC.
@@ 3.1.3 Intrinsic limitation of MIA as a Privacy Auditing Framework
MIAs are intrinsically incapable of identifying points close to training points, regardless of how similar they are, because these points are, by definition, non-members in the scope of MIAs. Hence, there is a huge space of points that contain private information but are deemed non-members in the current privacy auditing framework. In this way, MIAs as privacy auditing tools become bad when the queries move away from the original data. Figure 2 shows the MIAs under-perform on non-original data. This inspires our formulation of RaMIA, where these points will be classified as "members" for better and more comprehensive privacy auditing.
@@
Figure 2: MIA performance gets worse when the query points become further away from the training points. We define points different from training points but carry similar information as members. In 2(a), $m$ is the number of missing values. The query is constructed by filling in them with the most likely values. In 2(b), the point query changes to photos of the same identity who has at least one photo in the training set. In 2(c), the queries are vertically flipped images. In 2(d), $d$ is the Hamming distance to original sentences.
@@ 3.2.1 Range membership inference game
The main difference between the two games is that the adversary now receives a range query (Step 4 in Def 2) instead of a point query (Step 3 in Def 1). We assume that the adversary is able to sample within the range. This is a reasonable assumption because the adversary is usually assumed to have the ability to sample from the original data distribution $\pi$ [39, 44, 47]. Given a range query and a sampler of $\pi$, it is not difficult to sample within the range.
@@
A range can be defined by a center, which is a point, a radius representing the size of the range, and a distance function which the radius is defined with. We refer to the center as the query center, the radius as the range size, and the distance function as the range function in this paper. One way to visualize a range is to imagine a unit $l_{2}$ ball around a point $x$, replacing the radius and $l_{2}$ distance with any arbitrary choice of range sizes and functions. Our framework can cater to any arbitrary range function. It can be spatial based (e.g. $l_{p}$ distances), transformation based (e.g. geometric transformations) and semantic based (e.g. owner/main features of the data). In the experiment section, we will present results with all of these types of range functions. Note that our attack reduces to user-level inference [31, 23, 27, 11, 10] when the range function is user-based.
@@
In Step 3 of the range membership inference game, the details of how the challenger samples the ranges are intentionally omitted. This is because the ranges can be constructed around either in-distribution or out-of-distribution data points for both in- and out-ranges. The details of how we construct the in- and out-ranges for our experiments are elaborated in Appendix B.
@@ 3.3 Evaluation of RaMIA
Similarly, we evaluate RaMIA with AUCs. However, the notion of true positives and false positives are different from those defined in MIA, as both are defined on the range level. To avoid confusion, we call them Range TPR and Range FPR, which means a range is correctly/wrongly predicted to contain at least one training point.
@@ 4.1 Composite hypothesis testing
Similar to likelihood ratio tests for membership inference games A.1, we can also construct two hypotheses for range membership inference game (Def 2):
@@
To make full use of the Bayes Factor, we need to know the prior distribution, which is unrealistic. On the other hand, taking the max seems to be a more intuitive approach for range membership inference attacks, because it provides a two-step solution: search and test. Searching for the points with the highest membership score is conceptually equivalent to identifying the points that are most likely to be training points. However, this assumes that we can reliably find the max values in a given range. Since most ranges are large data subspaces, it is very challenging to find the optimal points within the large space. Even if the search space can be navigated, any search algorithm is likely to return local maxima. Hence, a robust way is to aggregate the top samples. However, membership inference attacks are known to be unreliable on out-of-distribution (OOD) data [47] and can assign them high scores. When the sampling space contains only these data as opposed to real and in-distribution (ID) data, the maximum might not be anything close to true training data, increasing the FPR and lowering the AUC as a result.

### v2

@@ Abstract
Machine learning models can leak private information about their training data, but the standard methods to measure this risk, based on membership inference attacks (MIAs), have a major limitation. They only check if a given data point exactly matches a training point, neglecting the potential of similar or partially overlapping data revealing the same private information. To address this issue, we introduce the class of range membership inference attacks (RaMIAs), testing if the model was trained on any data in a specified range (defined based on the semantics of privacy). We formulate the RaMIAs game and design a principled statistical test for its composite hypotheses. We show that RaMIAs can capture privacy loss more accurately and comprehensively than MIAs on various types of data, such as tabular, image, and language. RaMIA paves the way for a more comprehensive and meaningful privacy auditing of machine learning algorithms.
@@ 1 Introduction
To quantify the privacy risk of machine learning models, a privacy notion needs to be agreed upon first. The reigning privacy notion is defined by membership information. Membership information of a data point is binary, but this single bit of information carries huge privacy implications. Being able to infer membership information opens up the possibility of conducting data reconstruction attack [37, 17, 4, 29], where the reconstruction attack inspects the membership of plausible data points to recover the training set. The de-facto way to audit the privacy risk according to this privacy notion is to conduct membership inference attacks (MIAs) [40], where an adversary aims to predict whether a given query data is part of the training set of the target model. The more powerful the membership inference attack is, the higher the privacy risk the target model bears.
@@
Membership inference attacks provide a lower bound of the true privacy risk of a model, so improving the attack performance also means tightening the bound of privacy risk estimation. So far, the community has put much effort into improving the power of membership inference attacks by crafting better membership signals and constructing better statistical tests [39, 40, 36, 45, 5, 48]. While these have been useful for the betterment of privacy auditing, they have ignored the fundamental drawback of membership inference attacks as a practical privacy auditing tool, i.e., MIAs assume it is a privacy concern only if the adversary can identify the exact, full version of the training data. However, if the adversary can identify data points that are similar enough to the training points, it should also be treated as a significant privacy risk, because those points can contain similar levels of private information. For example, two Alice’s photos taken from slightly different angles, or with a different background would contain similar private information about Alice’s face or location. Similar to geometric transformations in images, small perturbations or rephrasing in textual data also affect little in the sensitivity of the information conveyed [13]. This oversight means private information leakage beyond exact matches of training data is currently ignored and the contemporary privacy auditing tools might produce overly optimistic results.
@@
Besides, focusing on exact membership inference attacks renders them incapable of handling queries with missing values. This is another major limitation of MIAs, as data records with the same sensitive features and a few missing non-private features could carry a similar level of private information as the full data records. Imagine the case where an adversary can infer that an Asian person of age 25, identification number 123456 is in an HIV patients dataset, there is no need to identify or collect the exact values of the rest of the features, because the adversary is already able to pinpoint who has HIV from the given subset of key features. Even if the key identifiers are missing and unknown, it is still a grave privacy threat if the attacker can infer the rest of the features from the subset of known features that often contain quasi-identifiers, which has been extensively studied in the failure of k-anonymity as a privacy-preserving measure [38], where the attacker can reconstruct and identify training data with certain features masked. With existing inference attacks, there are two types of approaches to work with incomplete data. The first approach is to impute the missing features before performing membership inference attacks. However, if we simply make up the missing values or pass noisy data to the membership inference attack, the attack is expected to output “not a member,” since the chance of us coming up with the right values is extremely slim. If we use attribute inference [20] to help fill in the missing features, the imputed data may have inflated membership scores even if they are non-members, leading to a higher false positive rate. Therefore, existing MIA-based privacy auditing frameworks does not provide a natural solution to quantifying privacy leakage with missing features. The second approach that works with incomplete data is property inference [41, 1, 8, 9, 46], but it aims to extract statistical information about the features instead of testing membership information of data records. Hence, none of the existing inference attack frameworks can reliably quantify privacy leakage with incomplete data entries.
@@
We argue that privacy quantification should not be point-based, because a small neighborhood around training points, defined based on a distance function out of the semantics of privacy, also contains similar private information. Hence, in this paper, we are proposing a new inference attack framework called range membership inference attacks (RaMIAs) to better capture the notion of privacy. Instead of using point queries and testing for exact matches, range membership inference attacks use range queries that cover a set of points. The goal of range membership inference attacks is to infer if the given range query contains any training point. In this paper, we denote the “neighborhood” as the “range”, and its corresponding distance function as the “range function”.
@@
Range membership inference attacks extend the formulation of membership inference attacks. We adapt the original inference game formulation to reflect the change to range queries in RaMIAs. This extended formulation produces composite hypotheses in the likelihood ratio tests, which are the standard and best attack techniques in MIAs [39, 45, 5, 48]. Our method is based on standard statistical methods for composite hypothesis testing, namely generalized likelihood ratio tests (GLRTs) and Bayes factors. We show that RaMIAs can provide a more comprehensive notion of privacy by detecting private information leakage from the vicinity of training data when MIAs underestimate such privacy risk. Specifically, we observe a simple flipping can cause the membership score to decrease from very high to 0 (Figure 1), and the overall AUC can drop 20% if we test image classifiers with horizontally flipped images (Figure 2(c)). This is actually the expected behavior of MIAs since the transformed training data are non-members by definition. RaMIA, implemented with our simple attack algorithm (Sec 4), surpasses MIA by at least 5% on image datasets (Fig 3(b), 3(c)), providing better privacy auditing at the cost of as few as 15 samples, which is insignificant compared to the dimensionality of the data space.
@@
In this paper, we emphasize the motivation and formulation of our newly proposed attack framework, RaMIA. As a proof-of-concept, we implement RaMIA with a simple attack strategy and experiment on tabular, image and text datasets, where RaMIA unanimously outperforms MIA. Additionally, our attack can also be potentially used in the pioneering membership inference and data extraction attacks on generative models [4, 44, 7], where the current evaluation requires finding the closest training image for all candidate data and computing their distances. By setting the distance function as the range function and conducting RaMIAs, we can evaluate the attacks more systematically.
@@
Figure 1: Examples of similar data with drastically different MIA scores. MIA would classify one of them as member, but the other as non-member. The dog image is from the CIFAR-10 dataset, while the text data is from AG News. The image classifer is not trained with horizontal flipping augmentation.
@@ 2.1 Membership inference attacks
The membership inference attack (MIA) [40] is a type of inference attack against machine learning models to infer whether a given data sample is part of the model’s training set. Mathematically, given a model $f$ and a query point $\mathbf{x}$, the MIA aims to output 1 if $\mathbf{x}$ is a training point, and 0 otherwise. There are various methods to construct and conduct the attack, and it is still an active research direction that sees more powerful attacks being developed. Shokri et al. [40] use a shadow model-based approach where shadow models are trained on known training sets in similar ways to the target model. Confidence values of the training and test data on the shadow models are computed, which are then used as benchmarks in testing. However, the high cost and strong assumption of knowing the target model’s training details make the attack often infeasible. Yeom et al. [46] use model loss as a signal and threshold it, scraping the need for shadow models. Then MIA is formulated as an inference game (See Sec 3.1.1). Researchers turn to the principled approach to solve the game via likelihood ratio tests [39, 5, 45, 48]. Carlini et al. [5] and Ye et al. [45] propose reference-model based approaches, where target signals are compared to those obtained on reference models to obtain the likelihood ratio. To further boost the attack power, Zarifzadeh et al. [48] assumes the attacker has access to a pool of population data so that the likelihood ratio from reference-based attacks can be calibrated on ratios obtained on (non-member) population data.
Membership inference attacks with augmentations
Recent attacks [5, 48] find extra attack performance on image data from augmenting the test queries with train-time augmentations, as these augmented training images could have been seen by the model during training. However, this assumes that the attacker knows the exact train-time augmentations in advance, and is able to sample from them. Augmenting training images with non train-time augmentations is not considered and allowed for valid reasons: those augmented images would be non-members in the current privacy notion based on membership inference, and testing them as members would violate the current privacy notion.
@@ 2.2 Range queries
If we make a connection to the field of databases, the existing membership inference attack operates on point queries or exact match queries. That is, each query to the membership inference attack only contains one data point and the attack only concerns if this very point is in the training set. On the other hand, range query, which is also a common querying operation in database systems, wants to retrieve all data points that fall into the ”range”. The most fundamental difference to point query is that the retrieved result often contains multiple data points instead of a single one. Our proposed attack, the range membership inference attack, operates with range queries.
@@ 3.1.2 Evaluation of MIA
Evaluation is done with a set of training and test points. True positive rate (TPR) and false positive rate (FPR) are computed by sweeping over all possible threshold values. By plotting the receiver operating characteristic curve (ROC), the power of an attack strategy can be represented by the area under the curve (AUC). A clueless adversary who can only randomly guess the membership labels is expected to get an AUC of 0.5. For stronger adversaries, they predict membership more accurately at each error level. Hence, they would achieve higher TPR at each FPR, and get a higher AUC.
@@ 3.1.3 Intrinsic limitation of MIA as a Privacy Auditing Framework
MIAs are intrinsically incapable of identifying points close to training points, regardless of how similar they are, because these points are, by definition, non-members in the scope of MIAs. Hence, there is a huge space of points that contain private information but are deemed non-members in the current privacy auditing framework. In this way, MIAs as privacy auditing tools become bad when the queries move away from the original data. Figure 2 shows the MIAs under-perform on non-original data. This inspires our formulation of RaMIA, where these points will be classified as ”members” for better and more comprehensive privacy auditing.
@@
Figure 2: MIA performance gets worse when the query points become further away from the training points. We define points different from training points but carry similar information as members. In 2(a), $m$ is the number of missing values. The query is constructed by filling in them with the most likely values. In 2(b), the point query changes to photos of the same identity who has at least one photo in the training set. In 2(c), the queries are horizontally flipped images. In 2(d), $d$ is the Hamming distance to original sentences.
@@ 3.2.1 Range membership inference game
The main difference between the two games is that the adversary now receives a range query (Step 4 in Def 2) instead of a point query (Step 3 in Def 1). We assume that the adversary is able to sample a set of points covered by any given range, which effectively means sampling within the range. This is a reasonable assumption because the adversary is usually assumed to have the ability to sample from the original data distribution $\pi$ [40, 45, 48] as specified in the membership inference game formulation. Given a range query and a sampler of $\pi$, it is not difficult to sample within the range.
@@
A range can be defined by a center, which is a point, a radius representing the size of the range, and a distance function which the radius is defined with. We refer to the center as the query center, the radius as the range size, and the distance function as the range function in this paper. One way to visualize a range is to imagine a unit $l_{2}$ ball around a point $x$, replacing the radius and $l_{2}$ distance with any arbitrary choice of range sizes and functions. Our framework can cater to any arbitrary range function that is potentially privacy-preserving. It can be spatially based (e.g. $l_{p}$ distances), transformation based (e.g. geometric transformations), and semantic based (e.g. owner/main features of the data). In the experiment section, we will present results with all of these types of range functions. Note that our attack reduces to user-level inference [31, 23, 27, 11, 10] when the range function is user-based, which means the privacy concern is whether any other data from the same owner are used in training. We will comment on the relations between RaMIA and other existing inference attacks later in the paper.
@@
In Step 3 of the range membership inference game, the details of how the challenger samples the ranges are intentionally omitted. This is because the ranges can be constructed around either in-distribution or out-of-distribution data points for both IN and OUT-ranges. The details of how we construct the iIN and OUT-ranges for our experiments are elaborated in Section 5.1.
@@ 3.3 Evaluation of RaMIA
Similar to MIAs, we evaluate RaMIA with AUCs. However, the notion of true positives and false positives are different from those defined in MIA, as both are defined on the range level. Specifically, if a range contains at least one training point, it is IN; otherwise, it is OUT. The TPR in RaMIA is the proportion of IN-ranges that are correctly identified by the attacker. On the other hand, the FPR in RaMIA is the proportion of OUT-ranges that yield positive predictions by the attacker. To avoid confusion, we call them (Range) TPR and (Range) FPR.
@@ 4 Range membership inference attacks
4.1 (Simple) Hypothesis testing
$H_{0}$ $:\text{The given }z\text{ is not a training point }(b=0).$
$H_{1}$ $:\text{The given }z\text{ is a training point }(b=1).$
The likelihood ratio test (LRT) is then conducted
$\frac{\text{I\kern-1.49994ptP}{\left(\theta|H_{1}\right)}}{\text{I\kern-1.49994ptP}{\left(\theta|H_{0}\right)}}$ (1)
This is usually called ”simple” hypothesis testing because each $H$ contains a single hypothesis. The scoring function in the membership inference attack can be considered as an approximation of the likelihood function $\text{I\kern-1.49994ptP}{(\cdot)}$.
@@ 4.2 Composite hypothesis testing
Similar to likelihood ratio tests for membership inference games 4.1, we can also construct two hypotheses for range membership inference game (Def 2):
@@
[…]
Since it can be intractable to iterate over all points in a given range $\mathcal{R}_{b}$, we use a set of sampled points $S$ in the given range as its proxy. In this way, the two hypotheses are:
[…]
The likelihood ratio in this case is still $\frac{\text{I\kern-1.04996ptP}{\left(\theta|H_{1}\right)}}{\text{I\kern-1.04996ptP}{\left(\theta|H_{0}\right)}}$. However, the alternative hypothesis $H_{1}$ is composite because it is a union of multiple hypotheses $\bigcup_{z_{i}\leftarrow S}(z_{i}\in D)$. Therefore, we need to use statistical methods tailored for composite hypothesis testing. There are two commonly used methods for it: Bayes Factor [22] and Generalized Likelihood Ratio Tests (GLRTs) [43].
GLRT
GLRT assumes that the true hypothesis $h^{*}$ is explicitly listed in the composite hypothesis, which, in our case, means the training point is in the set of samples $\{z_{i}\}$. In this case, the composite hypothesis can be reduced to a single hypothesis by taking the maximum:
$\text{I\kern-1.49994ptP}{\left(\theta|H_{1}\right)}\Rightarrow\max_{x\in S}\text{I\kern-1.49994ptP}{(\theta|x\in D)}.$ (2)
Bayes Factor
Bayes Factor assumes the hypothesis is a random variable, each with a likelihood value according to a prior distribution. In our case, it means each point $z$ in the range can be sampled with probability determined by the prior data distribution $\text{I\kern-1.49994ptP}{(z)}$. To reduce the composite hypothesis to a simple one, Bayes Factor computes the weighted average:
[…]
This can be thought of as taking the expectation of the composite hypothesis or using the most representative hypothesis instead of the maximal (as in GLRT) in the likelihood ratio test.
Why both methods fall short in RaMIA
@@
At first look, GLRT might be an intuitive choice, because taking the maximum leads to a two-step strategy: search and test. Searching for the points with the highest membership score is conceptually equivalent to identifying the points that are most likely to be training points, and their membership should be indicative of the ranges’ membership. However, this assumes the true training point is included in every sampled set created by the adversary, which is extremely unlikely. Secondly, this also assumes that we can reliably find the max values in a given range. Since most ranges are large data subspaces, it is very challenging to find the optimal points within the large space. Even if the search space can be navigated, any search algorithm is likely to return local maxima.
To make full use of the Bayes Factor, we need to know the prior distribution, which is also not possible.
A more important issue is that the power of existing statistical methods relies on the absolute correctness of the likelihood values: the larger the likelihood values, the more likely the point is a member. In other words, the largest likelihood values of non-members must be smaller than the smallest likelihood of members. However, membership inference attacks are known to be non-perfect; they are also known to be particularly unreliable on out-of-distribution (OOD) data [48] and can assign them high scores. When the sampling space contains only these data as opposed to real and in-distribution (ID) data, the maximum might not be anything close to true training data, increasing the FPR and lowering the AUC as a result.
4.3 Our approach: Trimmed averages